The greenhouse I am building for my wife is 6 feet longer than it is wide. I only have enough concrete for 56ft of foundation (the number of feet in the foundation is equal to the perimeter of the building) what are the possible values for the length of the building?
Answer by Earlsdon(6103)

Since you have specified the length in terms of the width, you can write:
L = W+6
The perimeter is 56 ft:
P = 2L + 2W Replace the L here with W+6 and solve for W.
56 = 2(W+6)+2W
56 = 2W+12+2W
56 = 4W+12 Subtract 12 from both sides.
44 = 4W Divide both sides by 4.
11 = W
The width is 11 feet and the length is:
L = W+6 = 11+6 = 17 feet
Check:
Perimeter = 2L+2W = 2(17) + 2(11) = 34 + 22 = 56 feet.